OpenStudy (anonymous):

Which Value(s) of X are excluded from the solution to the following equation ? Show all of your work to get to your solution . x^2-5x-20 x 2 ----------------- = -------- + -------- x^2-17x+70 x-10 x-7

OpenStudy (aravindg):

factorise the numerator and denominator

OpenStudy (anonymous):

How do you do that ?

OpenStudy (anonymous):

x ------ Would X = 10 ? x-10

OpenStudy (aravindg):

yep !

OpenStudy (aravindg):

similarily for 2/x-7

OpenStudy (aravindg):

at these points the fraction becomes undefined

OpenStudy (anonymous):

So im going to have 3 diffrent answers ?

OpenStudy (aravindg):

3? hwy 3 ?

OpenStudy (anonymous):

Well Thats a NO ! But im still a little confused . SO 2 ------------- x= 9 ? x-7

OpenStudy (aravindg):

equate denominator to 0 x-7=0 x=?

OpenStudy (anonymous):

how did u get a 0 ?

OpenStudy (aravindg):

because denomiator cannot be 0 for a fraction I considered that situation

OpenStudy (anonymous):

oh okay . So x=7

OpenStudy (aravindg):

yep :)

OpenStudy (anonymous):

x^2-5x-20 ---------------- = x = ?? x^2-17x+70

OpenStudy (aravindg):

the conditions x=7,9 are to be excuded from answer..I think that answers ur question

OpenStudy (anonymous):

I dont get it . Whats x ?

OpenStudy (aravindg):

x=7 and 9 !

OpenStudy (anonymous):

Oh okay . But i have to show work . And i dont know how u got that answer .

OpenStudy (aravindg):

just by equation denomiators to 0 x-10 =0 x=10 x-7=0 x=7 oops its 10 not 9 :)

OpenStudy (aravindg):

*equating

OpenStudy (anonymous):

Okay but what about the first part of the problem. x^2-5x-20 ---------------- How do i find x here ? x^2-17x+70

OpenStudy (anonymous):

x^2-17x+70 factors to (x-7)(x-10) so you already have eliminated the x-values that don't work.